A wire 2.60 m long with a cross-sectional area of 7.07x10-6 m2 is connected to a 30.0 V power supply. If a current of 12.0 A is measured in the wire,

a) what is the resistivity of the wire in ohm-meters? Answer in scientific notation.

b) If this resistivity had been measured at 20.0 °C, what temperature, in Celsius, would we need to heat/cool the wire to for the resistance to be 4.25 Ω if the temperature coefficient of resistivity is 5.25x10-3/°C.

a) ______ Ω m

b) ______ °C
Answer #1

PL (@) R D R = >/H B. 30 12 = 2.5 2 I e R = Pt 2.6 2.5 = 7.0710 -6 -6 Pa 6.798 X 10 2-mb P Sot axt T , gan so tx T 12 (P. - f.) = a *(Ti - T2) © R₂ = 4.252 =) Sex H A 7.078 106 (1 4.25 2.6 11.5578106 X Se -3 @ 5.